Find the equation of the line with a slope of 1- containing g the point (-6, -2)
Answer by Alan3354(69443)

Find the equation of the line with a slope of 1- containing g the point (-6, -2)
---------------
Use y-y1 = m*(x-x1) where m = slope and (x1,y1) is the point.